# SOAP: Enhancing Spatio-Temporal Relation and Motion Information Capturing for Few-Shot Action Recognition

### v4

@@ 3.1. Problem Formulation
Classifying an unlabeled query into one of the support classes is the inference goal of FSAR. The support set has at least one labeled sample per class. Following the prior works (cao2020few; perrett2021temporal; wang2022hybrid; wang2023molo), we randomly select few-shot tasks from training set for episodic training. In each task, the support set $S$ contains NN classes and $K$ samples per class, i.e., NN-way $K$-shot setting. The query set $Q$ in each task has samples to be classified. Uniformly sampling $F$ frames of each video, we define the $k^{\text{th}}\left(k=1,\cdots,K\right)$ sample of the $c^{\text{th}}\left(c=1,\cdots,N\right)$ class of support set $S$ as:
@@ 3.2. Overall Architecture
$3.$ 3-Dimension Enhancement Module
@@ 4.1. Experimental Configuration
In 3DEM, the size of $\mathrm{Conv3D}$ is $3\times 3\times 3$. For CWEM, the expand channel number $C_{r}$ is set to 16, while $\mathrm{Conv1D}$ size is 33. We set $\mathcal{O}=\left\{1,2,3\right\}$ in HMEM, the $\mathrm{Conv2D}^{M}$ with shared parameters are also $3\times 3$ in each branch. All 2D convolution for changing channel number is set to $1\times 1$. Except for these channel recovery convolutions, other convolutions maintain the same shape of input and output. We employ widely-used ResNet-50 (he2016deep) or ViT-B (dosovitskiy2020image) as the backbone initialized with pre-trained weights on ImageNet (deng2009imagenet). The final outputs of the backbone are 2048-dimensional vectors ($D=2048$). For three linear layers, the parameters are randomly initialized while $d_{k}$ and $d_{v}$ are both set to 1152.
@@ 4.3. Essential Components and Factors
$4.$ Research on How SOAP Works
@@ Appendix D Hyperparameter Study
Table IV illustrates the SOAP-Net structure in our implementation, as detailed in the main paper. The $\mathrm{Conv3D}$ size is $3\times 3\times 3$ in 3DEM. In CWEM, the expand channel number $C_{r}$ is set to 16, while $\mathrm{Conv1D}$ size is 33. $\mathcal{O}=\left\{1,2,3\right\}$ and the $\mathrm{Conv2D}^{M}$ sizes are $3\times 3$ in HMEM. For a supplementary interpretation of implementation details, extensive experiments are conducted on individual modules with diverse configuration to determine the optimal hyperparameters.
@@ Acknowledgements.
The authors would like to appreciate all participants of peer review. Wenbo Huang sincerely thanks Bingxiao Shi (USTB) and all family members for the encouragement at an extremely difficult time (from 2022.9 to 2023.9). This work is supported by the National Key Research and Development Program for the 14th-Five-Year Plan of China under Grants 2023YFC3804104, 2023YFC3804100, National Natural Science Foundation of China under Grants No.62072099, 62232004, 62373194, 62276063, Jiangsu Provincial Key Laboratory of Network and Information Security under Grant No.BM2003201, Key Laboratory of Computer Network and Information Integration (Ministry of Education, China) under Grant No.93K-9, and the Fundamental Research Funds for the Central Universities.

### v5

@@ 3.1. Problem Formulation
Classifying an unlabeled query into one of the support classes is the inference goal of FSAR. The support set has at least one labeled sample per class. Following the prior works (cao2020few; perrett2021temporal; wang2022hybrid; wang2023molo), we randomly select few-shot tasks from training set for episodic training. In each task, the support set $S$ contains $N$ classes and $K$ samples per class, i.e., $N$-way $K$-shot setting. The query set $Q$ in each task has samples to be classified. Uniformly sampling $F$ frames of each video, we define the $k^{\text{th}}\left(k=1,\cdots,K\right)$ sample of the $c^{\text{th}}\left(c=1,\cdots,N\right)$ class of support set $S$ as:
@@ 3.2. Overall Architecture
3.3. 3-Dimension Enhancement Module
@@ 4.1. Experimental Configuration
In 3DEM, the size of $\mathrm{Conv3D}$ is $3\times 3\times 3$. For CWEM, the expand channel number $C_{r}$ is set to 16, while $\mathrm{Conv1D}$ size is $3$. We set $\mathcal{O}=\left\{1,2,3\right\}$ in HMEM, the $\mathrm{Conv2D}^{M}$ with shared parameters are also $3\times 3$ in each branch. All 2D convolution for changing channel number is set to $1\times 1$. Except for these channel recovery convolutions, other convolutions maintain the same shape of input and output. We employ widely-used ResNet-50 (he2016deep) or ViT-B (dosovitskiy2020image) as the backbone initialized with pre-trained weights on ImageNet (deng2009imagenet). The final outputs of the backbone are 2048-dimensional vectors ($D=2048$). For three linear layers, the parameters are randomly initialized while $d_{k}$ and $d_{v}$ are both set to 1152.
@@ 4.3. Essential Components and Factors
4.4. Research on How SOAP Works
@@ Appendix D Hyperparameter Study
Table IV illustrates the SOAP-Net structure in our implementation, as detailed in the main paper. The $\mathrm{Conv3D}$ size is $3\times 3\times 3$ in 3DEM. In CWEM, the expand channel number $C_{r}$ is set to 16, while $\mathrm{Conv1D}$ size is $3$. $\mathcal{O}=\left\{1,2,3\right\}$ and the $\mathrm{Conv2D}^{M}$ sizes are $3\times 3$ in HMEM. For a supplementary interpretation of implementation details, extensive experiments are conducted on individual modules with diverse configuration to determine the optimal hyperparameters.
@@ Acknowledgements.
The authors would like to appreciate all participants of peer review. Wenbo Huang sincerely thanks Bingxiao Shi (USTB) and all family members for the encouragement at an extremely difficult time. This work is supported by the National Key Research and Development Program for the 14th-Five-Year Plan of China 2023YFC3804104 in 2023YFC3804100, National Natural Science Foundation of China under Grants No.62072099, 62232004, 62373194, 62276063, Jiangsu Provincial Key Laboratory of Network and Information Security under Grant BM2003201, Key Laboratory of Computer Network and Information Integration of MOE China under Grant No.93K-9, and the Fundamental Research Funds for the Central Universities.